### v3

# Functional vs. Object-Oriented: Comparing How Programming Paradigms Affect the Architectural Characteristics of Systems

ABSTRACT.
This study compares the impact of adopting object-oriented programming (OOP) or functional programming (FP) on the architectural characteristics of software systems. For that, it examines the design and implementation of a Digital Wallet system developed in Kotlin (for OOP) and Scala (for FP). The comparison is made through a mixed-method approach. The self-ethnographic qualitative analysis provides a side-by-side comparison of both implementations, revealing the perspective of those writing such code. The survey-based quantitative analysis gathers feedback from developers with diverse backgrounds, showing their impressions of those reading this code. Hopefully, these results may be useful for developers seeking to decide which paradigm is best suited for their next project.
Software Architecture, Programming Paradigms, Object-Oriented Programming, Functional Programming, Empirical Software Engineering
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Brazilian Symposium on Software Quality; Nov 04–07, 2025; São José dos Campos, SPcopyright: none
1. Introduction
After decades of dominance of Object-Oriented Programming (OOP), the Functional Programming (FP) paradigm has gained significant attention in the software industry (Finley, 2022). This shift is a response to growing complexities in software systems, the demand for better scalability, and the need for more predictable, robust, and maintainable codebases.
According to the IEEE Spectrum magazine in 2022 (Scalfani, 2022), one of the clearest indications of this shift is the increasing incorporation of functional features into different mainstream programming languages. Evaluating how FP compares to OOP is essential to understanding the future of software development.
The main research question of this paper can be summarized as:
RQ: How do the functional and object-oriented paradigms impact different architectural characteristics of a system?
According to Richards and Ford (Richards and Ford, 2020), an architectural characteristic describes a concern that follows three criteria:
  1. (1)
    it specifies a non-domain design consideration,
  2. (2)
    it influences structural aspects of the design, and
  3. (3)
    it is critical or important to the application’s success.
This research focuses on six architectural characteristics, summarized in table 1. They encompass concerns that highlight important differences between the paradigms, making it easier to compare their strengths and limitations.
Table 1. Evaluation Criteria for Qualitative Analysis
Architectural Characteristic From the code, how easy it is to…
Extensibility …add new behavior
Reusability …apply it in a new use case
Error Handling …understand errors handled by it
Error Propagation …understand errors propagated by it
Testability …test its intended behavior
Readability …understand its intended behavior
Due to space constraints, this paper will address only four of the six architectural characteristics from table 1, whose analyses brought the most interesting results. For a full discussion, please refer to the bachelor thesis that originated this paper (Dias et al., 2024).
sRQ1: How do the functional and object-oriented paradigms impact the extensibility of a system?
Section 1 aims to explore how each paradigm influences the ability to add new components to the system. OOP typically relies on object composition and class inheritance, while FP uses high-order functions and generic data structures.
sRQ2: How do the functional and object-oriented paradigms impact the reusability of a system?
Section 1 aims to explore how each paradigm influences the ability to reuse components in the system. OOP typically relies on design patterns that take advantage of polymorphism, while FP combines list-based immutable data types with closures and recursion.
sRQ3: How do the functional and object-oriented paradigms impact error handling and propagation of a system?
Section 1 aims to explore how each paradigm embodies distinct philosophies in error management. OOP typically uses exceptions, while FP applies monadic types.
2. Programming Paradigms
Understanding the principles of OOP and FP is essential for a meaningful comparison between them. Their distinct philosophies, strengths, and limitations significantly influence how software systems approach architectural characteristics.
2.1. Object-Oriented Programming
According to Martin (Martin, 2017), any OOP language must support these three characteristics: encapsulation, inheritance, and polymorphism.
Encapsulation
It ensures that a clear boundary is established around a cohesive set of data and behaviors (Martin, 2017). Within this boundary, data is hidden and protected from direct external access, while only specific functions are exposed to interact with the data. This approach is commonly seen in object-oriented programming through private data members and public methods in classes.
Inheritance
It is the process of reusing and extending a group of variables and functions defined in one scope within another (Martin, 2017). In OOP, this often means that a new class (called a subclass) can inherit properties and behaviors from an existing class (called a superclass). This allows the new class to reuse, override, or expand existing functionality without redefining it entirely.
Polymorphism
It is the ability of a method to behave differently based on the context in which it is used (Martin, 2017). Polymorphism unblocks dependency inversion by allowing high-level modules to depend on abstractions rather than concrete implementations. It decouples implementation details from the modules that depend on them through interfaces, allowing dependencies to be reversed.
2.2. Functional Programming
According to Pilquist et al. (Pilquist et al., 2023), functional programming is based on the premise of writing programs using only pure functions, which leads to other patterns like high-order functions and monadic types.
Pure Functions
They are functions without side effects (Martin, 2017). A function has side effects if it does something other than returning a result, such as modifying a variable, throwing an exception, or printing to the console.
High-Order Functions
In FP, functions are values (Pilquist et al., 2023). They can be assigned to variables, stored in data structures, and passed as arguments to functions. A function that accepts other functions as arguments is called a higher-order function (HOF).
Monadic Types
Most FP languages support pattern matching (Pilquist et al., 2023) as a way to execute functions depending on the value passed to them. Constructs such as Option and Either have the properties of Monads (Pilquist et al., 2023), types that leverage pattern matching to provide clean composition and error handling.
3. Related Literature
Few works in the literature focus on the same research questions proposed in section 1. A search with the keywords “Functional versus Object-Oriented Programming” on on Google Scholar results in papers that propose to mix the paradigms in the context of a single programming language or application.
In the late 1990s, Harrison et al. (Harrison et al., 1996; Samaraweera and Harrison, 1998) compared the code of $12$ programs produced in SML (FP) and C++ (OOP), measuring their difference on development and code metrics. In 2016, Alic et al. (Alic et al., 2016) compared algorithms in Java and C# (OOP) versus Haskell and F# (FP), focusing on lines of code and efficiency. In 2025, Khan et al. (Khan et al., 2025) compared the impact of choosing OOP, FP or Procedural Programming for implementing microservices architectures (Newman, 2021), analyzing their impact on application design.
The current gap in the literature is an opportunity to explore the differences between paradigms once again, using modern programming languages and focusing on both code and design.
4. Methodology
This section details the research methodology employed in this project, whose goal is to answer section 1.
4.1. Proof of Concept
The first phase of the research consists of the specification and development of a proof-of-concept system that can be used as an object of study. This provides a detailed example for comparing the object-oriented and functional paradigms.
A Digital Wallet was chosen as the proof-of-concept system for this research. It contains many functionalities, such as wallet management, transaction management, and calculation of balances. These operations demand intricate data handling to keep consistency, whose implementation can be approached differently within the object-oriented and functional paradigms.
To compare the paradigms, the same Digital Wallet system was developed twice, following the good practices of each paradigm.
Object-Oriented PoC
The object-oriented version of the Digital Wallet system was based on the Kotlin programming language, which supports the OOP principles from section 2.1. The developer can define classes in a straightforward manner, enabling the structuring of data and behavior through objects.
Functional PoC
The functional version of the Digital Wallet system was based on the Scala programming language, which supports classic FP principles from section 2.2. The developer can focus on immutability and pure functional constructs, aligning with the core principles of functional programming.
Section 5 describes the set of functional requirements that guided the development of both versions of the Digital Wallet system. Sections 6 and 7 describe the main business logic that was implemented in both proofs of concept. The code developed during the research is available in the paper’s reproduction package.
4.2. Analysis
The second phase of the research consists of an analysis of the paradigms using a mixed-method approach. By triangulating benefits and drawbacks of each paradigm over the architectural characteristics, it is possible to provide an answer to section 1.
The analysis will be based on the implementation of the Digital Wallet system. In particular, the focus is on code snippets that perform the same task from both proofs of concept.
To provide a good perspective over the paradigms, this research proposes to combine two research techniques:
Self-Ethnographic Qualitative Analysis
It represents the perspective of writing code in the different paradigms. In particular, the self-ethnographic methodology means that it presents the assessments of the authors of the paper, who developed the proofs of concept. These results are further explored in section 8.
Survey-Based Quantitative Analysis
It represents the perspective of reading code in the different paradigms. In particular, the survey-based methodology means that it summarizes the assessments of different developers with varying levels of experience with the paradigms. These results are further explored in section 9.
After presenting these results, section 10 combines these perspectives to answer the research subquestions.
5. Requirements
The Digital Wallet system is the object of study used in this work to establish the comparison between the object-oriented and the functional programming paradigms.
Although it models a real-world problem, the Digital Wallet system was not developed for production but rather was designed as a proof of concept whose functionalities highlight the differences between the paradigms. Therefore, the Digital Wallet system ignores some important non-functional requirements for this type of system, such as security and auditability.
The functional requirements outlined for the Digital Wallet system are listed below.
5.1. Wallet Management
The Digital Wallet system should provide each customer with three different types of wallets:
Real Money Wallet
It represents a real-world checking account owned by the customer. This wallet is the entry point of the Digital Wallet system, where the customer can deposit from or withdraw to an external bank account. It is modeled to be a temporary allocation of the customer’s funds until they are moved to another wallet.
Investment Wallet
It represents a portfolio of investment options where customers can allocate their funds. Funds are invested by transferring funds from the Real Money Wallet to the Investment Wallet and are liquidated the other way around. Investments and liquidations can be initiated at any time but will only settle on the next business day.
Emergency Funds Wallet
It represents deposit insurance where customers can allocate funds without taking the risks of the Investment Wallet. Funds are deposited into and withdrawn from the Emergency Funds Wallet through instant transfers to and from the Real Money Wallet.
5.2. Transaction Lifecycle Management
Figure 1. Transaction Lifecycle. :
The Digital Wallet system should support the creation, validation, and execution of transactions between wallets. Transactions represent the flow of money inside the Digital Wallet system, which might succeed or fail.
To provide customers with all the functionalities, the Digital Wallet system needs to support the following transactions:
  • Deposit, which represents a deposit to the Real Money Wallet, modeling the entry of funds into the Digital Wallet system.
  • Withdrawal, which represents a withdrawal from the Real Money Wallet, modeling the exit of funds from the Digital Wallet system.
  • Transfer, which represents an instant transfer to or from the Real Money Wallet. While the Digital Wallet system is not required to handle transfers directly with real-world bank accounts, it should integrate with a third-party API to enable this functionality.
  • Hold, which represents an amount of money reserved for future use. This transaction ensures funds until an application settles.
  • Transfer From Hold, which represents an instant release of an amount of money previously reserved in a Hold. This transaction settles an application, using the available funds.
Figure 1 summarizes the compatibility between transactions and different wallets. Every transaction must be validated before execution. Deposit and Withdrawal transactions are restricted to the Real Money Wallet. Transfer transactions are allowed only between Real Money Wallet and Emergency Funds Wallet. Hold transactions can be placed only on Real Money Wallet and Investment Wallet, whereas Transfer From Hold transactions are limited to being executed between them. Finally, all transactions – except for Deposit – require a balance check to ensure sufficient funds are available to complete the operation.
The Digital Wallet system must manage transaction failures as part of the transaction lifecycle. Transactions failing validation should be marked as permanently failed, while those that fail during execution should be eligible for retry. In general, validation should capture expected, unrecoverable errors, such as insufficient funds. Execution errors, though unlikely after successful validation, may still occur due to internal issues or third-party API instabilities. In such cases, the system should allow to retry the transaction.
5.3. Investment Customization
The Digital Wallet system should allow customizing investments and liquidations. The participation should be a percentage of the total invested under an Investment Wallet, and the customer should have real-time control over it. This requirement includes the capacity of setting the percentage to zero for undesired investments or even selecting a single investment.
5.4. Minor Requirements
Financial Tracking
The Digital Wallet system must track all financial transactions to maintain accurate account balances and ensure data integrity.
Investment and Liquidation Settlement
The Digital Wallet system must settle all investment and liquidation requests initiated on the previous business day.
Batch Processing of Transactions
The Digital Wallet system must support batch processing of transactions with atomicity. In any given batch, all transactions must either complete successfully or, if any transaction fails, the entire batch must be rolled back. This ensures data integrity and prevents partial updates that could disrupt the financial record-keeping system.
6. Digital Wallet System: Entities
This section presents entities and value objects (Evans, 2003) representing the core business concepts of the Digital Wallet system.
6.1. Wallet
The Wallet entity models an account where customer funds are held and managed. It functions as the hub for all financial transactions, allowing money to be deposited or withdrawn to other destinations. The Digital Wallet system provides three types of wallets, as specified in section 5.1Real Money Wallet, Investment Wallet, and Emergency Fund Wallet.
While Transaction entities dictate the flow of funds between Wallets entities, the latter are unaware of the former’s existence. Their role is to model and hold customer funds within the system. They are not responsible for determining the validity of any incoming or outgoing Transaction. This separation of concerns ensures that Wallet entities focus on balance management without handling Transaction logic.
6.2. Subwallet
The Subwallet concept is introduced to represent the distribution of funds within the same wallet.
The Real Money Wallet has only one Subwallet entity: the Real Money Subwallet. Similarly, the Emergency Funds Wallet entity has only one Subwallet entity: the Emergency Funds Subwallet. In practice, having a single Subwallet means that, for record-keeping, funds in the Wallet are not divided in any way.
Figure 2. Wallets and corresponding Subwallets. :
The relevance of the Subwallet entity becomes evident in the Investment Wallet, where funds are allocated across various investment options, as illustrated by figure 2. Introducing a new investment option involves composing new Subwallet entities within the Investment Wallet.
6.3. Transaction
A Transaction entity represents an attempt to change the state of a Wallet. Each transaction specifies an amount, as well as originator and beneficiary, which can either be a Wallet instance (and its designated Subwallet) or an external bank account. Each Transaction also has a status representing its current phase in processing. Possible statuses include Processing, Failed, Transient Error, and Completed.
Transaction Types
A Transaction can be classified as a Deposit, Withdrawal, Hold, Transfer, or Transfer From Hold. Each transaction type represents a standard operation that modifies the state of a Wallet as specified in section 5.2.
Transaction validation
A Transaction must be validated to ensure they adhere to the internal rules of the Digital Wallet system. For example, a Deposit and Withdrawal can only be initiated on a Real Money Wallet, but not on an Investment Wallet. These validations follow the restrictions specified in section 5.2.
Transaction status
A Transaction is created with the Processing status, indicating they have not yet gone through validation or execution. The Failed status is a terminal state, representing a permanent failure with no possibility of retry. Similarly, the Completed status is a terminal state, representing that all validations and execution steps succeeded, the transaction has settled, and the involved Wallet entity was updated. The Transient Error status is a cyclic state, representing a recoverable error that occurred during transaction processing (such as a failure in a third-party API call), which is eligible for retry.
6.4. Journal Entry
A Journal Entry entity documents the effects of a Transaction on a Wallet. It tracks changes in the Wallet balance. By listing all Journal Entry entities linked to a Wallet, it is possible to calculate its current state accurately.
Journal Entry entities are always created in pairs to represent the source and destination of funds. To accurately reflect changes in a Wallet, a Journal Entry includes the following:
  • wallet_id: It identifies the source or target Wallet. It may be absent to represent an external bank account.
  • subwallet_id: It identifies the Subwallet under a Wallet. It may also be absent to represent an external bank account.
  • amount: It accounts for the money being added (if positive) or deducted (if negative) from the Subwallet.
  • balance_type: It can have three values: Internal, for funds sent to or received from an external bank account; Available, for funds added to or deducted from the available balance of a Subwallet; and Holding, for funds added to or deducted from the reserved balance of a Subwallet.
The Ledger entity collects all Journal Entry entity created by the Digital Wallet system. Each Transaction with Completed status results in a pair of Journal Entry being recorded in the Ledger. Figure 3 illustrates how these journal entries look like for each transaction type.
Figure 3. Pairs of Journal Entry for each Transaction type. :
6.5. Investment Policy
An Investment Policy entity defines the allocation of an investment portfolio by associating each Subwallet with a specific percentage representing its share of the total portfolio, as specified in section 5.3.
Only a Subwallet within the Investment Wallet is permitted to be included in an Investment Policy.
7. Digital Wallet System: Services
The Digital Wallet system relies on a set of services that collectively operate the functionalities described in section 5.
Figure 4 summarizes the existing services and illustrates the communication flow between them.
Figure 4. Services that compose the Digital Wallet system. :
7.1. Ledger Service
The Ledger service is responsible for orchestrating between other services and the Digital Wallet system Ledger entity. It provides two functionalities: (1) posting new Journal Entry, and (2) querying the Ledger for its balance. For the latter, it aggregates Journal Entry to compute the balance of a specific Wallet.
7.2. Investment Service
The Investment service is dedicated to handling investment and liquidation requests, which involve multiple steps and do not settle instantly like other transaction types. This service knows how to effectively use an Investment Policy to initiate and manage the transactions required to complete an investment or liquidation.
7.3. Transaction Service
The Transactions service provides functionalities to fulfill each step of a transaction lifecycle specified in section 5.2. It creates, validates, and processes Transactions, encapsulating the business logic for each Transaction type.
This service also guarantees that the Ledger is accurately updated upon Transaction completion, maintaining consistency across the system.
Although the Transactions service is responsible for managing different Transaction types, it remains agnostic to their specific purposes. Its role is to ensure that all transactions are processed correctly and that the ledger remains consistent.
7.4. Wallet Service
The Wallets service is responsible for managing and initiating Transaction s within Wallet s. It encapsulates the business logic needed to translate money movements initiated by the customer into a series of Transaction s that fulfill these requests.
The types of requests handled by the Wallets service include:
  • Deposit: Adds funds into the Real Money Wallet.
  • Withdraw: Removes funds from the Real Money Wallet.
  • Emergency Allocation: Adds funds into the Emergency Funds Wallet.
  • Emergency Release: Transfers funds from the Emergency Funds Wallet to the Real Money Wallet.
  • Investment: Transfers funds from the Real Money Wallet to the Investment Wallet, distributing the amount across investment options based on the InvestmentPolicy.
  • Liquidation: Transfers funds from the Investment Wallet to the Real Money Wallet, liquidating investments proportionally according to the InvestmentPolicy.
8. Qualitative Analysis
The qualitative analysis compares functionally equivalent code snippets from the Kotlin and Scala proofs of concept, evaluating them according to the architectural characteristics focused on by the research sub-questions. This methodology is self-ethnographic, since it presents the assessments of the authors who led the development process.
8.1. Extensibility [section 1]
To analyze the Extensibility quality attribute, this section examines the Kotlin and Scala implementations of a Wallet in the Digital Wallet system.
Kotlin
[…]
List of codes 1 Wallet class inheritance hierarchy in Kotlin :
In Kotlin, the Wallet entity is structured using OOP principles like inheritance and polymorphism. The abstract class Wallet serves as a blueprint for all specific wallet types: Real Money Wallet, Investment Wallet, and Emergency Funds Wallet. This is illustrated in listing 1.
In Kotlin, getAvailableBalance is an abstract method of the Wallet class, which encapsulates the logic for querying balances from the ledger. All subclasses must implement their specific logic for computing balances. This ensures consistency across wallet types while allowing each subclass to customize its behavior. Such an approach leverages polymorphism: the behavior of a method depends on the subclass of the object being called.
A future requirement demanding a new Wallet could be achieved by simply adding a new subclass that inherits from the Wallet abstract class. This subclass automatically gains access to the shared attributes defined in the superclass while also being required to implement its abstract method, enabling it to define specific balance-query behavior.
Scala
In Scala, the Wallet entity is structured using FP principles like immutable data structures. The case class Wallet consolidates all possible attributes and contains an additional field WalletType that distinguishes the wallet derivations within the system. This is illustrated in listing 2.
[…]
List of codes 2 Wallet type and factory pure function in Scala :
In Scala, getAvailableBalance is a standalone function in the WalletService class, which encapsulates the logic for querying balances from the ledger. For every variant of the WalletType enumeration, the function specifies a corresponding list of ledger queries to compute the available balance. This provides behavior tailored to each wallet type without requiring separate case classes. Such an approach leverages pattern matching: the walletType field maps to different balance computation logic.
A future requirement demanding a new Wallet could be achieved by simply adding a new WalletType enumeration. In Scala, pattern matching is exhaustive, meaning that any unhandled enumeration variant results in a compilation error. Therefore, pattern matching ensures that all standalone functions related to WalletType need to be revisited, ensuring it to define all balance-query behavior.
Kotlin vs. Scala
Both Kotlin and Scala implementations promote extensibility, enabling the system to adapt through localized code updates. In Kotlin, this is achieved by adding a new subclass to represent a wallet type and implementing the specialized getAvailableBalance method for that class. In Scala, extensibility is achieved by introducing a new variant to the WalletType enumeration.
These features also enforce the extension of existing specialized logic whenever a new wallet type is introduced. While Kotlin requires the creation of a dedicated class and method implementation for each new type, in Scala this can be achieved by simply adding a new enumeration variant and updating the existing functions that centralize the behavior of all wallet types.
8.2. Reusability [section 1]
To analyze the Reusability quality attribute, this section examines the Kotlin and Scala implementations of retrying a batch of transactions. Not all transaction failures within the Digital Wallet system are permanent. For example, transactions that pass validation but fail during execution are eligible for retry.
The retryBatch method attempts to reprocess each transaction within a batch up to n times. If any transaction fails while interacting with the partner API, the entire retryBatch operation should fail. Conversely, if all transactions are successfully processed, the batch should be marked as successful.
Kotlin
[…]
[ numbers=left, label=( Kotlin — retryBatch ), ]code/kotlin/retryBatch.kt
List of codes 3 retryBatch built with a for loop in Kotlin
In Kotlin, the implementation of the retry logic is built around a while loop. This is illustrated in listing 3. It leverages the counter variable attempts to track the number of retries for each transaction and the flag variable success to indicate whether the transaction processing succeeded.
This design lacks the flexibility to be reused for other operations, even though it provides the necessary control for transactions.
Scala
[…]
List of codes 4 retryBatch built over the HOF retry in Scala
In Scala, the implementation of the retry logic is built around a higher-order function named retry. This is illustrated in listing 4. It takes a function f: => Either[A, B] and an integer n as parameters to execute f up to n times, early returning Right if any attempt succeeds, or returning Left if all n attempts fail.
This design encapsulates the retry logic into a reusable function, providing this functionality to any function that needs to be retried.
Kotlin vs. Scala
This example demonstrates how the concept of higher-order functions enables the reuse of the same logic across multiple operations in the Scala PoC, eliminating the need to replicate similar logic for different functions. By leveraging pattern matching and recursion, the Scala PoC also achieves a more concise and expressive code structure compared to the Kotlin PoC.
8.3. Error Handling and Propagation [section 1]
To analyze error handling and propagation, this section examines how Kotlin and Scala implement the logic to create investments.
An investment allocates funds from the Real Money Wallet entity to the Investment Wallet entity. This operation consists essentially of a hold Transaction that does not settle instantly.
Kotlin
[…]
[ numbers=left, label=( Kotlin — invest ), ]code/kotlin/invest.kt
List of codes 5 invest method using exceptions in Kotlin
In Kotlin, the hold Transaction is created through the processTransaction method, which can throw exceptions. These exceptions must be explicitly caught in the invest method to map them to a corresponding WalletService exception. Additionally, within the catch blocks, it is necessary to ensure the Transaction transitions to the appropriate state. This is illustrated in listing 5.
However, if a new exception type is created in the future, developers must explicitly update the try-catch block, adding a new catch case, otherwise risking unhandled runtime exceptions if they forget about it.
Additionally, as a try-catch block might not handle all the error cases, reasoning about the program becomes more challenging since the code does not clearly express the potential failure paths.
Scala
[…]
[ numbers=left, label=( Scala — invest ), ]code/scala/invest.sc
List of codes 6 invest function using monadic types in Scala
In Scala, errors are emitted via the monadic type Either. This approach allows them to be explicitly modeled as part of the method’s return type, making it clear to developers that a method can fail. This is illustrated in listing 6.
The use of Either enhances type safety by enforcing that the caller must handle both success and failure cases, ensuring that no error is missed during development. This promotes more predictable and reliable error handling within the system.
9. Quantitative Analysis
The quantitative analysis aims to gather feedback from developers with diverse backgrounds through a survey, complementing the qualitative analysis.
The draft survey was initially reviewed by a developer with experience in both paradigms. Then, it was beta-tested with two additional developers knowledgeable in both paradigms. After incorporating all feedback, the final version was made publicly available to a wider public.
The survey was divided into two sections:
Background
This section was designed to map the developers’ experience with object-oriented and functional programming languages, since such experience might influence participants’ reasoning when evaluating the code snippets.
Code Analysis
This section was designed to measure the developer’s impressions about the two Digital Wallet system proof of concepts. Each question includes a brief explanation of a functionality from the Digital Wallet system, followed by two code snippets: the first from the Kotlin PoC and the second from the Scala PoC. For each code snippet, participants were asked to evaluate the six architectural characteristics summarized in table 1. They answered the question: “From the code, how easy is it to…” using a Likert scale with five steps, from Very Easy to Very Hard.
At the publication of the bachelor thesis that inspired this paper (Dias et al., 2024), the survey received eight responses. Considering space constraints, this section will only address two of four questions proposed in the questionnaire, which intersect with some of the code previously analyzed in section 8. For a full analysis, please refer to the bachelor thesis (Dias et al., 2024).
The full questionnaire and corresponding responses can be found in the paper’s reproduction package.
9.1. Question #1 — retryBatch
Question #1 collected feedback on the implementation of the retryBatch, whose code was explored in listings 3 and 4.
The results of Question #1 are shown in figure 5. They reveal a notable difference in the architectural characteristics of error handling, error propagation, and readability.
Many developers found the Scala implementation of the retry logic to be more challenging to handle, to propagate errors, and to read when compared to the Kotlin implementation. This indicates that, for these aspects, Kotlin may offer a clearer approach.
Figure 5. Kotlin vs. Scala — retryBatch :
9.2. Question #3 — invest
Question #3 collected feedback on the implementation of the invest logic, whose code was explored in listings 5 and 6.
The results of the Question #3 are shown in figure 6. They reveal that extensibility is slightly favored in the Scala PoC compared to the Kotlin PoC. On the other hand, testability showed a significant advantage in favor of Kotlin.
Error handling and error propagation were comparable between both implementations, though some responses suggested these aspects might be more challenging to achieve in Scala. Reusability was also quite similar for both implementations, with a slight preference for Scala. Readability, while inconclusive for Scala, received more favorable feedback for Kotlin.
Figure 6. Kotlin vs. Scala — invest:
10. Discussion
Based on the findings from the qualitative and quantitative analyses, this section discusses the research subquestions.
sRQ1: How do the functional and object-oriented paradigms impact the extensibility of a system?
The qualitative analysis revealed that both Kotlin and Scala offer features that promote extensibility. Kotlin leveraged object-oriented principles such as inheritance and polymorphism to achieve extensibility, while Scala employed enumerations and pattern matching to accomplish the same objective. Although the Scala implementation was more concise, requiring less boilerplate code to extend a model, both approaches were effective and provided comparable benefits to the extensibility characteristic.
The quantitative analysis corroborated these findings. The perceptions of Scala and Kotlin implementations were similar, with only minor differences observed. Notably, these differences mostly favored the Kotlin implementation of the Digital Wallet System.
sRQ2: How do the functional and object-oriented paradigms impact the reusability of a system?
Scala demonstrated stronger support for code reusability through higher-order functions, enabling generic implementations of various functionalities. In contrast, the Kotlin implementations in both examples lacked sufficient flexibility to be effectively adapted for different domains.
However, the quantitative analysis did not reveal significant differences between Kotlin and Scala in terms of reusability. Responses indicated that both implementations were perceived as similar, with only minor variations.
The discrepancy between the qualitative and quantitative analysis findings suggests that the reusability characteristic may considerably depend on the specific functionality being evaluated.
sRQ3: How do the functional and object-oriented paradigms impact error handling and propagation of a system?
Kotlin propagates errors using exceptions and handles them with traditional try-catch blocks. In contrast, Scala utilizes the monadic type Either[A, B] to propagate errors. Error handling in Scala is performed through higher-order functions, allowing mapping to new error types as needed.
By incorporating errors directly into return types, the Scala implementation enforces comprehensive error handling through type checking, ensuring that new error cases cannot be missed.
Nevertheless, the quantitative analysis did not reflect the same dominance of the Scala implementation. Most responses indicated that Kotlin is either comparable or slightly superior to Scala in terms of error handling and propagation.
11. Threats to Validity
The threats to validity acknowledge the limitations of this research.
Construct Validity
The evaluation of programming paradigms in this research may be influenced by the choice of programming languages representing each paradigm. Kotlin and Scala were chosen as modern, multi-platform languages that run on the Java Virtual Machine (JVM), ensuring a level of syntactic compatibility between the programming languages. Despite Kotlin supporting FP features and Scala supporting OOP features, this research intentionally limited both to their selected domains.
External Validity
The qualitative analysis conducted in this research may reflect biases influenced by the author’s background and expertise. To address this, a mixed-methods approach was employed, incorporating quantitative analysis to complement the qualitative findings. Nevertheless, due to time constraints, the number of participants in the survey was limited. To strengthen the results, more participants should be included to triangulate the conclusions from this research.
12. Conclusion
RQ: How do the functional and object-oriented paradigms impact different architectural characteristics of a system?
This paper compared the impact of adopting OOP versus FP over six architectural characteristics of software systems. It introduced the Digital Wallet system, implemented twice: once in Kotlin (OOP), once in Scala (FP). The use of these modern, popular programming languages for the research will hopefully reach the interest of practitioners, who may understand the results and then be better informed when choosing a certain paradigm.
For future work, this research can be extended to target other architectural characteristics, such as scalability and security. Additionally, the current results can be enhanced by expanding the survey, thus addressing some of the threats to validity.